OpenStudy (bekkah323):

2x/x^2-16 plus x/x-4 *equation below

OpenStudy (bekkah323):

\[\frac{ 2x }{ x^2-16 }+\frac{ x }{ x-4 }\]

OpenStudy (anonymous):

do u want to simplify it?

OpenStudy (anonymous):

Firstly, factorise the bottom of the first fraction using the difference of two squares :)

OpenStudy (anonymous):

Then put both fractions under the same denominator (x+4)(x-4).

OpenStudy (bekkah323):

so r u saying to first factor the first denominator?

OpenStudy (anonymous):

yes :) First do that

OpenStudy (bekkah323):

ok so would it be\[\frac{ x^2+6x }{ (x-4) (x+4) }\]

OpenStudy (anonymous):

Yes, now multiply both the numerator and denominator of the second fraction such that both fractions have the same denominator :)